U throw a stone horizontally at a speed of 2.9m/s from the top of cliff that is 37 m high. How long does it take to reach the bottom . How far from the base does the stone hit the ground?
1
Expert's answer
2020-04-17T10:47:55-0400
"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(37)}{9.8}}=2.75\ s"

"x=vt=(2.9)(2.75)=8.0\ m"